OpenStudy (anonymous):

which fraction is the reciprocal of 4 1/9 A.36/1 B.37/9 C.9/4 D.9/37

OpenStudy (anonymous):

please help me

OpenStudy (mertsj):

Multiply them and find out. Reciprocals are two numbers whose product is 1

OpenStudy (anonymous):

The whole number 1 can be written as 9/9. So 4 is 4 x 1 = 4(9/9) = 36/9. Now 4 1/9 = 36/9 + 1/9 = 37/9. The reciprocal of this is 9/37, hence the answer is d.

OpenStudy (anonymous):

because 4 and 1/9 is the same as 37/9 and the reciprocal of that (flipped) is 9/37

OpenStudy (anonymous):

first of all....do you know how to change a mixed number to an improper fraction ?